How long will R60,000 accumulate to R62,430 if the simple interest rate agreed is 1.50%?

1
Expert's answer
2021-11-21T15:34:59-0500

Interest "=62,430-60,000=2,430"


"SI=\frac {PRT} {100}"


"2430=\frac {60000×1.5×T} {100}"


"243000=90,000T"


"T=\frac {243000} {90000}=2.7"


"Time =2.7yrs"